OpenStudy (haleyelizabeth2017):

What are the real or imaginary solutions of the polynomial equation?

OpenStudy (haleyelizabeth2017):

\(64x^3+27=0\)

OpenStudy (anonymous):

largely depends on the book

OpenStudy (alexandervonhumboldt2):

for first x=3/4

OpenStudy (haleyelizabeth2017):

okay :) I know how to solve the second one when I plug it into the quadratic formula. Thank you! I just needed help getting started :)

OpenStudy (alexandervonhumboldt2):

yeah!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (haleyelizabeth2017):

lol what's got you all excited?? :P

OpenStudy (alexandervonhumboldt2):

sorry the first () should be (4x+3) so x=-3/4

OpenStudy (haleyelizabeth2017):

Okay :) thanks!

OpenStudy (alexandervonhumboldt2):

(4x+3)*(16x^2-12x+9)

OpenStudy (haleyelizabeth2017):

Thank you again :)

OpenStudy (haleyelizabeth2017):

Does this look set up correctly @AlexandervonHumboldt2?\[x=\frac{ -(-12) \pm \sqrt {(-12)^2-4(16)(9)}}{ 2(16) }\]

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (haleyelizabeth2017):

Okay, great :)

OpenStudy (haleyelizabeth2017):

How can I simplify \(\sqrt {432}*i\)?